# Finite-Sample Non-Parametric Bounds with an Application to the Causal Effect of Workforce Gender Diversity on Firm Performance

### v1

@@ B.4 Proof of Corollary 1
$\Pr\left(\sup_{y}\left[F_{n}(y)-F(y)\right]>t\right)\leq\exp\left(-2nt^{2}\right),$
@@
ensures $\Pr(\mathcal{E}_{5}^{c})=\Pr(\mathcal{E}_{6}^{c})\leq\alpha_{u}/6$. The four mean- and share-events $\mathcal{E}_{1}$$\mathcal{E}_{4}$ and their bounds are unchanged, hence each still receives probability $\alpha_{u}/6$. Because the six complements jointly spend at most $\alpha_{u}$, Boole’s inequality and the algebra in Proposition 2 give
$\Pr\left(\Re_{u}\in H_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha_{u},\qquad u=m_{0},\dots,m_{1}.$
B.5 Proof of Proposition 3
Similar to the proof of Proposition 2, define the six “good” events $\mathcal{E}_{1}\cdots,\mathcal{E}_{6}$ and apply Boole’s (union) bound. We then choose each threshold $t_{i}$ so that $\Pr(\mathcal{E}_{i}^{c})\leq\alpha_{u}/6$ for $u=m_{0},\cdots,m_{1}$ under dependence:
@@ B.5 Proof of Proposition 3
$\Pr\left(\lvert\hat{\mu}_{k}-\mu_{k}\rvert\geq t_{k}\right)\leq T_{1}(t_{k})+T_{2}(t_{k})+T_{3}(t_{k}),$
@@
$T_{1}(t)$ $=N_{k}\exp\left(-\frac{(N_{k}\,t)^{\gamma}}{C_{1}}\right),$ (term 1)
$T_{2}(t)$ $=\exp\left(-\frac{(N_{k}\,t)^{2}}{C_{2}(1+N_{k}V)}\right),$ (term 2)
$T_{3}(t)$ $=\exp\left(-\frac{(N_{k}\,t)^{2}}{C_{3}\,N_{k}}\exp\left(\frac{(N_{k}\,t)^{\gamma(1-\gamma)}}{C_{4}(\log(N_{k}t))^{\gamma}}\right)\right).$ (term 3)
To make each term $\leq\alpha_{u}/18$ for $u=m_{0},\cdots,m_{1}$:
1. $T_{1}(t)\leq\alpha_{u}/18$ iff
$\frac{(N_{k}t)^{\gamma}}{C_{1}}\geq\log\frac{18N_{k}}{\alpha_{u}}\quad\Longrightarrow\quad t\geq t_{k}^{(1)}:=\frac{\left(C_{1}\log(18N_{k}/\alpha_{u})\right)^{1/\gamma}}{N_{k}}.$
2. $T_{2}(t)\leq\alpha_{u}/18$ iff
$\frac{(N_{k}t)^{2}}{C_{2}(1+N_{k}V)}\geq\log\frac{18}{\alpha_{u}}\quad\Longrightarrow\quad t\geq t_{k}^{(2)}:=\frac{\sqrt{C_{2}(1+N_{k}V)\log(18/\alpha_{u})}}{N_{k}}.$
3. $T_{3}(t)\leq\alpha_{u}/18$ iff
$\frac{(N_{k}t)^{2}}{C_{3}N_{k}}\exp\left(\frac{(N_{k}t)^{\gamma(1-\gamma)}}{C_{4}(\log(N_{k}t))^{\gamma}}\right)\geq\log\frac{18}{\alpha_{u}},$
so define $t_{k}^{(3)}$ to be the unique positive solution of this equation. Finally set
$t_{k}=\max\{t_{k}^{(1)},t_{k}^{(2)},t_{k}^{(3)}\},$
@@ 3. $T_{3}(t)\leq\alpha_{u}/18$ iff
Solving $2\exp(-A)\leq\alpha_{u}/6$ with $A=N_{k}\,t^{2}/[2(1+4C_{\upalpha})^{2}]$ gives
@@
so $\Pr(\mathcal{E}_{3}^{c})=\Pr(\mathcal{E}_{4}^{c})=\alpha_{u}/6$.
@@
By Lemma 5,
$\Pr\left(\sup_{y}\bigg{\lvert}F_{N_{k}}^{(k)}(y)-F^{(k)}(y)\bigg{\rvert}>t_{k}\right)\leq 2\exp\left(-\frac{N_{k},t_{k}^{2}}{2(1+4C_{\upalpha})^{2}}\right).$
Similarly, set
@@
so $\Pr(\mathcal{E}_{5}^{c})=\Pr(\mathcal{E}_{6}^{c})=\alpha_{u}/6$.
Combining these gives
$\Pr\left(\bigcap_{i=1}^{6}\mathcal{E}_{i}\right)\geq 1-\sum_{i=1}^{6}\Pr(\mathcal{E}_{i}^{c})=1-\alpha_{u}.$
On $\bigcap_{i}\mathcal{E}_{i}$, elementary algebra shows $\Re_{u}\in H_{\alpha_{u}}[\Re_{u}]$, so
$\Pr\left(\forall u\in\mathcal{M},\Re_{u}\in H_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha.$
@@
B.6 Proof of Proposition 4
Let $(Y_{ijt},Z_{ijt})_{t=1}^{T}$ be strictly stationary and $\upalpha$–mixing with $C_{\upalpha}=\sum_{r\geq 1}\upalpha(r)^{1/2}<\infty$. To simultaneously obtain a the $100(1-\alpha)\%$ confidence set for both the upper and lower bounds for the identification region (48), we must first find the confidence bounds with an appropriate significance level and combine them using Bonferroni inequalities, such that the combined confidence set has a $(1-\alpha)\%$ coverage rate, or:
$\begin{split}\Pr\left(\forall u\in\mathcal{M},\left\{[\mathcal{L}(\hat{\mathbf{\theta}}),\mathcal{U}(\hat{\mathbf{\theta}})]\subseteq\mathbb{I}(\Re_{u})\right\}\cap\left\{\sup_{y}\bigg{\lvert}F_{N_{0}}^{(0)}(y)-F^{(0)}(y)\bigg{\rvert}\leq\epsilon_{0}\right\}\right.\\ \left.\qquad\;\;\cap\left\{\sup_{y}\bigg{\lvert}F_{N_{1}}^{(1)}(y)-F^{(1)}(y)\bigg{\rvert}\leq\epsilon_{1}\right\}\right)&\geq 1-\alpha,\end{split}$ (73)
where $\hat{\mathbf{\theta}}=(\hat{\delta}_{1},\hat{\delta}_{0},\hat{p}_{1},\hat{p}_{0})^{\top}$ is a $4\times 1$ vector of estimators and $\mathbb{I}(\Re_{u})$ is an interval $l(\Re_{u}),u(\Re_{u})]$. In other words, for $u=m_{0},\cdots,m_{1}$, we wish to obtain:
@@ B.6 Proof of Proposition 4
$\Pr\left(\sup_{y}\bigg{\lvert}F_{N_{k}}^{(k)}(y)-F^{(k)}(y)\bigg{\rvert}\leq\epsilon^{(k)}\right)\geq 1-\frac{\alpha_{u}}{4},\quad\text{for}\quad k=0,1.$
Let us define the events:
@@
$\mathcal{E}_{3}$ $=\left\{\sup_{y}\bigg{\lvert}F_{N_{1}}^{(1)}(y)-F^{(1)}(y)\bigg{\rvert}\leq\epsilon_{1}\right\}$
$\mathcal{E}_{4}$ $=\left\{\sup_{y}\bigg{\lvert}F_{N_{0}}^{(0)}(y)-F^{(0)}(y)\bigg{\rvert}\leq\epsilon_{0}\right\}$
We know from Boole’s inequality that:
$\begin{split}\Pr\left(\bigcap\limits_{i=1}^{4}\mathcal{E}_{i}\right)&\geq 1-\sum\limits_{i=1}^{4}\Pr\left(\mathcal{E}_{i}^{c}\right)\\ &\geq 1-4\frac{\alpha_{u}}{4}\\ &\geq 1-\alpha_{u}.\end{split}$ (74)
Thus,
[…]
Additionally, we must ensure that:
$\Pr\left(\sup_{y}\bigg{\lvert}F_{N_{k}}^{(k)}(y)-F^{(k)}(y)\bigg{\rvert}>\epsilon_{k}\right)$ $\geq 2\exp\left(-\frac{N_{k}\epsilon_{k}^{2}}{2(1+4C_{\upalpha})^{2}}\right)$
$=\frac{\alpha_{u}}{4}\quad\text{for}\quad k=0,1$
which holds when
$\epsilon_{k}:=(1+4C_{\upalpha})\sqrt{\frac{2\log(8/\alpha_{u})}{N_{k}}}.$ (76)
[…]
$\sqrt{N}(\hat{\theta}-\theta)\xrightarrow{D}N(0,\Omega_{\theta}).$ (77)
From the proof of Proposition 1 and given the sample size in treatment group $k\in\{0,1\}$ and by
$Y^{(k)}_{(1)}\leq\cdots\leq Y^{(k)}_{(N_{k})},$
we know that
$\nabla\mathcal{L}(\theta)=\left(p_{1},-p_{0},\delta_{1}-U_{\alpha_{u}}^{(0)},L_{\alpha_{u}}^{(1)}-\delta_{0}\right)^{\top}$ (78)
$\nabla\mathcal{U}(\theta)=\left(p_{1},-p_{0},\delta_{1}-L_{\alpha_{u}}^{(0)},U_{\alpha_{u}}^{(1)}-\delta_{0}\right)^{\top}$ (79)
where unlike in Proposition 1,
$L_{\alpha_{u}}^{(k)}:=Y_{(1)}^{(k)}-\epsilon_{k}\quad\text{and}\quad U_{\alpha_{u}}^{(k)}:=Y_{(N_{k})}^{(k)}+\epsilon_{k}$ (80)
where $\epsilon_{k}$ is defined in Eq. (76). Similarly, the covariance matrix of estimators $\hat{\mathbf{\theta}}$ is given explicitly by:
Therefore,
$\sqrt{N}\left(\mathcal{L}(\hat{\theta})-\mathcal{L}(\theta)\right)$ $\xrightarrow{D}N\left(0,\nabla\mathcal{L}(\theta)^{\top}\Omega\nabla\mathcal{L}(\theta)\right)$ (82)
$\sqrt{N}\left(\mathcal{U}(\hat{\theta})-\mathcal{U}(\theta)\right)$ $\xrightarrow{D}N\left(0,\nabla\mathcal{U}(\theta)^{\top}\Omega\nabla\mathcal{U}(\theta)\right)$ (83)

### v2

@@ B.5 Proof of Corollary 1
$\Pr\left(\sup_{y}\left[F_{n}(y)-F(y)\right]>t\right)\leq\exp\left(-2N_{k}t^{2}\right),\quad\forall t\geq\sqrt{\frac{\ln 2}{2N_{k}}}$
@@
ensures $\Pr(\mathcal{E}_{5}^{c})=\Pr(\mathcal{E}_{6}^{c})=\alpha_{u}/6$. The four mean- and share-events $\mathcal{E}_{1}$-$\mathcal{E}_{4}$ and their bounds are unchanged, hence each still receives probability $\alpha_{u}/6$. Because the six complements jointly spend at most $\alpha_{u}$, Boole’s inequality and the algebra in Proposition 2 give
$\Pr\left(\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha_{u},\qquad u=m_{0},\dots,m_{1}.$
B.6 Proof of Proposition 3
Suppose the collection $(Y_{ijt},Z_{ijt})_{i,j,t}$ is a strictly stationary $\upalpha$-mixing process in the sense of Definition 1, and that conditional on $Z_{ijt}(\tau_{u})=k$, the potential outcomes have sub-exponential tails with $\psi_{1}$-norm bounded by $M_{k}$. The proof follows the same structure as Proposition 2: define the six “good” events $\mathcal{E}_{1},\dots,\mathcal{E}_{6}$ and apply Boole’s inequality. We choose each threshold $t_{i}$ so that $\Pr(\mathcal{E}_{i}^{c})\leq\alpha_{u}/6$ for $u=m_{0},\dots,m_{1}$.
@@ B.6 Proof of Proposition 3
$\Pr\left(\lvert\hat{\mu}_{k}-\mu_{k}\rvert\geq t_{k}\right)\leq T_{1}(t_{k})+T_{2}(t_{k})+T_{3}(t_{k}),\qquad k=0,1,$
@@
$T_{1}(t)$ $=N_{k}\exp\!\left(-\frac{(N_{k}t)^{\gamma}}{C_{1}}\right),$
$T_{2}(t)$ $=\exp\!\left(-\frac{(N_{k}t)^{2}}{C_{2}(1+N_{k}V)}\right),$
$T_{3}(t)$ $=\exp\!\left(-\frac{(N_{k}t)^{2}}{C_{3}N_{k}}\exp\!\left(\frac{(N_{k}t)^{\gamma(1-\gamma)}}{C_{4}(\log(N_{k}t))^{\gamma}}\right)\right).$
To enforce $\Pr(\mathcal{E}_{k}^{c})\leq\alpha_{u}/6$, it suffices to make each term $\leq\alpha_{u}/18$:
  1. 1.
    $T_{1}(t)\leq\alpha_{u}/18$ iff
    $t\geq t_{k}^{(1)}:=\frac{\left(C_{1}\log(18N_{k}/\alpha_{u})\right)^{1/\gamma}}{N_{k}}.$
  2. 2.
    $T_{2}(t)\leq\alpha_{u}/18$ iff
    $t\geq t_{k}^{(2)}:=\frac{\sqrt{C_{2}(1+N_{k}V)\,\log(18/\alpha_{u})}}{N_{k}}.$
  3. 3.
    $T_{3}(t)\leq\alpha_{u}/18$ defines a unique positive root $t_{k}^{(3)}$, since the LHS is strictly increasing in $t$.
Set
$t_{k}:=\max\{t_{k}^{(1)},\,t_{k}^{(2)},\,t_{k}^{(3)}\}.$
@@
Solving $2\exp(-A)=\alpha_{u}/6$ with $A=N_{k}t_{k}^{2}/[2(1+4C_{\upalpha})^{2}]$ gives
@@
so that $\Pr(\mathcal{E}_{3}^{c})=\Pr(\mathcal{E}_{4}^{c})=\alpha_{u}/6$.
@@
Lemma 6 (dependent DKW) states:
$\Pr\left(\sup_{y}\lvert F_{N_{k}}^{(k)}(y)-F^{(k)}(y)\rvert>t_{k}\right)\leq 2\exp\!\left(-\frac{N_{k}t_{k}^{2}}{2(1+4C_{\upalpha})^{2}}\right).$
Thus we may choose
@@
giving $\Pr(\mathcal{E}_{5}^{c})=\Pr(\mathcal{E}_{6}^{c})=\alpha_{u}/6$.
Step 1 Concluded.
By summing the six error probabilities,
$\Pr\left(\bigcap_{i=1}^{6}\mathcal{E}_{i}\right)\geq 1-\sum_{i=1}^{6}\Pr(\mathcal{E}_{i}^{c})\geq 1-\alpha_{u}.$
Step 2. From the intersection event to coverage.
On $\bigcap_{i=1}^{6}\mathcal{E}_{i}$, the perturbed means, shares, and CDF quantiles satisfy exactly the inequalities required in Lemma 1. Because Manski’s bounds are monotone in all these arguments, it follows algebraically that
$\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\qquad\text{whenever}\qquad(\hat{\theta},Y)\in\bigcap_{i=1}^{6}\mathcal{E}_{i}.$
Therefore,
$\Pr\left(\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right)\geq\Pr\left(\bigcap_{i=1}^{6}\mathcal{E}_{i}\right)\geq 1-\alpha_{u},\qquad u=m_{0},\dots,m_{1}.$
Finally, by the Bonferroni allocation $\sum_{u}\alpha_{u}=\alpha$,
$\Pr\left(\forall u\in\mathcal{M},\;\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha.$
@@
B.7 Proof of Proposition 4
Let $(Y_{ijt},Z_{ijt})_{t=1}^{T}$ be strictly stationary and $\upalpha$-mixing with $C_{\upalpha}=\sum_{r\geq 1}\upalpha(r)^{1/2}<\infty$. To simultaneously obtain the $100(1-\alpha)\%$ confidence set for both the upper and lower bounds for the identification region (47), we first construct, for each $u\in\mathcal{M}$, a random interval $\mathcal{H}_{\alpha_{u}}[\Re_{u}]$ with marginal coverage $1-\alpha_{u}$, then combine across $u$ by Bonferroni so that
$\Pr\left(\forall u\in\mathcal{M},\;\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha,\qquad\sum_{u=m_{0}}^{m_{1}}\alpha_{u}=\alpha.$
For fixed $u\in\mathcal{M}$, we aim at
$\begin{split}\Pr\Big(&[\mathcal{L}(\hat{\mathbf{\theta}}),\mathcal{U}(\hat{\mathbf{\theta}})]\subseteq\mathbb{I}(\Re_{u})\;\cap\;\big\{\sup_{y}\lvert F_{N_{0}}^{(0)}(y)-F^{(0)}(y)\rvert\leq\epsilon_{0}\big\}\\ &\cap\;\big\{\sup_{y}\lvert F_{N_{1}}^{(1)}(y)-F^{(1)}(y)\rvert\leq\epsilon_{1}\big\}\Big)\;\geq\;1-\alpha_{u},\end{split}$ (72)
where $\hat{\mathbf{\theta}}=(\hat{\delta}_{1},\;\hat{\delta}_{0},\;\hat{p}_{1},\;\hat{p}_{0})^{\top}$ is a $4\times 1$ vector of estimators and $\mathbb{I}(\Re_{u})$ is an interval $l(\Re_{u}),u(\Re_{u})]$.
Define the events
$\mathcal{E}_{1}$ $=\{\mathcal{L}(\hat{\theta})\geq l(\Re_{u})\},\qquad\mathcal{E}_{2}=\{\mathcal{U}(\hat{\theta})\leq u(\Re_{u})\},$
$\mathcal{E}_{3}$ $=\left\{\sup_{y}\lvert F_{N_{1}}^{(1)}(y)-F^{(1)}(y)\rvert\leq\epsilon_{1}\right\},\qquad\mathcal{E}_{4}=\left\{\sup_{y}\lvert F_{N_{0}}^{(0)}(y)-F^{(0)}(y)\rvert\leq\epsilon_{0}\right\}.$
By Boole’s inequality,
$\begin{split}\Pr\left(\bigcap_{i=1}^{4}\mathcal{E}_{i}\right)&\geq 1-\sum_{i=1}^{4}\Pr(\mathcal{E}_{i}^{c}).\end{split}$ (73)
If we choose the four components so that $\Pr(\mathcal{E}_{i}^{c})\leq\alpha_{u}/4$ for $i=1,\dots,4$, then $\Pr(\cap_{i=1}^{4}\mathcal{E}_{i})\geq 1-\alpha_{u}$.
Under $\upalpha$-mixing with $C_{\upalpha}<\infty$, a mixing-adjusted DKW bound gives, for $k=0,1$,
$\Pr\left(\sup_{y}\lvert F^{(k)}_{N_{k}}(y)-F^{(k)}(y)\rvert>\epsilon_{k}\right)\leq 2\exp\left(-\frac{N_{k}\epsilon_{k}^{2}}{2(1+4C_{\upalpha})^{2}}\right).$
Thus choosing
$\epsilon_{k}:=(1+4C_{\upalpha})\sqrt{\frac{2\log(8/\alpha_{u})}{N_{k}}},\qquad k=0,1,$ (74)
ensures $\Pr(\mathcal{E}_{k+2}^{c})\leq\alpha_{u}/4$ for $k=0,1$.
On the event $\mathcal{E}_{k+2}$ the uniform control $\sup_{y}\lvert F^{(k)}_{N_{k}}(y)-F^{(k)}(y)\rvert\leq\epsilon_{k}$ implies the quantile-sandwich
$F^{(k),-1}(p-\epsilon_{k})\leq\hat{F}^{-1}_{N_{k}}(p)\leq F^{(k),-1}(p+\epsilon_{k}),\quad p\in[\epsilon_{k},1-\epsilon_{k}].$
Fix a deterministic
$r_{k}:=\min\{\epsilon_{k},1/2-\epsilon_{k}-1/N_{k}\}$
and define the empirical quantile endpoints
$L^{(k)}:=\hat{F}^{-1}_{N_{k}}(r_{k}+\epsilon_{k})=Y^{(k)}_{\left(\lceil(r_{k}+\epsilon_{k})N_{k}\rceil\right)},\qquad U^{(k)}:=\hat{F}^{-1}_{N_{k}}(1-r_{k}-\epsilon_{k})=Y^{(k)}_{\left(\lceil(1-r_{k}-\epsilon_{k})N_{k}\rceil\right)}.$ (75)
Then on $\mathcal{E}_{k+2}$ we have
$F^{(k),-1}(r_{k})\leq L^{(k)},\qquad U^{(k)}\leq F^{(k),-1}(1-r_{k}),$
so $L^{(k)},U^{(k)}]$ are conservative surrogates for the unknown population tail quantiles $\left[F^{(k),-1}(r_{k}),F^{(k),-1}(1-r_{k})\right]$. Manski’s bounds are monotone in the support endpoints, so using $L^{(k)},U^{(k)}$ in place of the unknown limits preserves coverage of $\Re_{u}$.
For the mean and proportion components, we construct one-sided bounds
$\Pr\big(\mathcal{L}(\hat{\theta})\geq l(\Re_{u})\big)\geq 1-\frac{\alpha_{u}}{4},\qquad\Pr\big(\mathcal{U}(\hat{\theta})\leq u(\Re_{u})\big)\geq 1-\frac{\alpha_{u}}{4},$
by applying a CLT and delta method to
$\hat{\theta}=(\hat{\delta}_{1},\;\hat{\delta}_{0},\;\hat{p}_{1},\;\hat{p}_{0})^{\top},\qquad\theta=(\delta_{1},\;\delta_{0},\;p_{1},\;p_{0})^{\top}.$
Under $\upalpha$-mixing, a suitable CLT yields
$\sqrt{N}(\hat{\theta}-\theta)\xrightarrow{D}N(0,\Omega_{\theta}),$ (76)
for some positive semidefinite $4\times 4$ matrix $\Omega_{\theta}$, which can be consistently estimated (for example) by a heteroskedasticity- and autocorrelation-consistent estimator (see neweywest1987).
From Proposition 1, treating the support endpoints $L^{(k)},U^{(k)}$ as fixed constants, the gradients of the lower and upper functionals with respect to $\theta$ are
$\nabla\mathcal{L}(\theta)$ $=\left(p_{1},\;-p_{0},\;\delta_{1}-U^{(0)},\;L^{(1)}-\delta_{0}\right)^{\top},$ (77)
$\nabla\mathcal{U}(\theta)$ $=\left(p_{1},\;-p_{0},\delta_{1}-L^{(0)},\;U^{(1)}-\delta_{0}\right)^{\top}.$ (78)
A first-order Taylor expansion gives
$\mathcal{L}(\hat{\theta})-\mathcal{L}(\theta)=\nabla\mathcal{L}(\theta)^{\top}(\hat{\theta}-\theta)+o_{p}(N^{-1/2}),$
@@ B.7 Proof of Proposition 4
$\sqrt{N}\left(\mathcal{L}(\hat{\theta})-\mathcal{L}(\theta)\right)$ $\xrightarrow{d}N\left(0,\nabla\mathcal{L}(\theta)^{\top}\Omega_{\theta}\nabla\mathcal{L}(\theta)\right),$ (79)
$\sqrt{N}\left(\mathcal{U}(\hat{\theta})-\mathcal{U}(\theta)\right)$ $\xrightarrow{d}N\left(0,\nabla\mathcal{U}(\theta)^{\top}\Omega_{\theta}\nabla\mathcal{U}(\theta)\right).$ (80)
Let
$\text{S.E.}\big(\mathcal{L}(\hat{\mathbf{\theta}})\big):=\sqrt{\frac{1}{N}\nabla\mathcal{L}(\hat{\theta})^{\top}\hat{\Omega}_{\theta}\nabla\mathcal{L}(\hat{\theta})},\qquad\text{S.E.}\big(\mathcal{U}(\hat{\mathbf{\theta}})\big):=\sqrt{\frac{1}{N}\nabla\mathcal{U}(\hat{\theta})^{\top}\hat{\Omega}_{\theta}\nabla\mathcal{U}(\hat{\theta})}$,
where $\hat{\Omega}_{\theta}$ is a consistent estimator of $\Omega_{\theta}$. Then the one-sided Wald inequalities
@@
hold with probability at least $1-\alpha_{u}/4$ each, asymptotically. Combining these four one-sided events with (73) yields
$\Pr\left([\mathcal{L}(\hat{\mathbf{\theta}}),\mathcal{U}(\hat{\mathbf{\theta}})]\subseteq\mathbb{I}(\Re_{u})\right)\geq 1-\alpha_{u}.$
Finally, by $\sum_{u=m_{0}}^{m_{1}}\alpha_{u}=\alpha$ and another application of Boole’s inequality over $u$, we obtain
$\Pr\left(\forall u\in\mathcal{M},\;\Re_{u}\in\mathcal{H}_{\alpha_{u}}[\Re_{u}]\right)\geq 1-\alpha,$
which establishes (72).